The interior angle of a polygon is 100 . The other interior angles are all equal to 110 . How many sides has the polygon ?
Answer by stanbon(75887)

The interior angle of a polygon is 100 . The other interior angles are all equal to 110 . How many sides has the polygon ?
--------
The sum of the exterior angles is 360 degrees.
The 100 degree interior angle has an exterior of 80 degree
Each of the 110 degree interior angles has an exterior of 70 degrees
-------
Let "n" be the number of sides, then there are n exterior angles.
----
Exterior Angle Equation::
80 + (n-1)70 = 360 degrees
70(n-1) = 280
n-1 = 4
n = 5 (# of sides)
--------------
Cheers,
Stan H.
------------